### v1

# A Simple and Effective Pruning Approach for Large Language Models

Abstract
As their size increases, Large Languages Models (LLMs) are natural candidates for network pruning methods: approaches that drop a subset of network weights while striving to preserve performance. Existing methods, however, require either retraining, which is rarely affordable for billion-scale LLMs, or solving a weight reconstruction problem reliant on second-order information, which may also be computationally expensive. In this paper, we introduce a novel, straightforward yet effective pruning method, termed Wanda (Pruning by Weights and activations), designed to induce sparsity in pretrained LLMs. Motivated by the recent observation of emergent large magnitude features in LLMs, our approach prune weights with the smallest magnitudes multiplied by the corresponding input activations, on a per-output basis. Notably, Wanda requires no retraining or weight update, and the pruned LLM can be used as is. We conduct a thorough evaluation of our method on LLaMA across various language benchmarks. Wanda significantly outperforms the established baseline of magnitude pruning and competes favorably against recent methods involving intensive weight update. Code is available at https://github.com/locuslab/wanda.
1 Introduction
Large language models Touvron et al. (2023); Brown et al. (2020); OpenAI (2023); Hoffmann et al. (2022); Zhang et al. (2022); Scao et al. (2022) have recently reshaped the field of NLP with their remarkable performance across a range of complex language benchmarks Bommarito and Katz (2022); Wei et al. (2022b); Bubeck et al. (2023); Wei et al. (2022a). However, these models, with their billions of parameters, usually require significant computational resources. To democratize LLMs, considerable efforts have been taken to mitigate their high computational cost. Many of the notable advancements to date have centered on model quantization, a process where parameters are quantized into lower bit-level representations. The fast pace of LLM quantization research Dettmers et al. (2022); Xiao et al. (2023); Frantar et al. (2023); Yao et al. (2022); Dettmers and Zettlemoyer (2022); Xiuying et al. (2022); Lin et al. (2023); Ahmadian et al. (2023); Dettmers et al. (2023) has led to substantial resource savings for these models llama.cpp ; Sheng et al. (2023).
Network pruning Hassibi et al. (1993); LeCun et al. (1989); Han et al. (2015); Liu et al. (2019); Jonathan and Michael (2019); Molchanov et al. (2017), on the other hand, shrinks network sizes by removing specific weights from the model – essentially setting them to zero. Along with quantization, it is often considered another popular approach for compressing neural networks. However, it has received relatively little focus in compressing LLMs. This seems to contradict the trend of model compression in the pre-LLM era, where both approaches have received large amounts of research effort. A quick review of existing pruning methods reveals a possible reason: they typically require retraining Han et al. (2015); Blalock et al. (2020), training from scratch Jonathan and Michael (2019); Renda et al. (2020); Louizos et al. (2018); Chen et al. (2020) or even an extensive iterative process Zhu and Gupta (2017); Gale et al. (2019); Li et al. (2020). The sheer computational resources required by LLMs limit these methods. A recent LLM pruning approach, SparseGPT Frantar and Alistarh (2023), does not require traditional retraining, but still demands a computationally intensive weight update process.
The argument concerning the need for retraining and weight update does not fully capture the challenges of pruning LLMs. One might reasonably expect to obtain a fairly high-performing initialization point for retraining using existing popular pruning methods. However, a recent study Frantar and Alistarh (2023) finds that magnitude pruning Han et al. (2015); Zhu and Gupta (2017); Gale et al. (2019), a well-established pruning approach, fails dramatically on LLMs with relatively low levels of sparsity. Considering the past success of magnitude pruning on smaller networks, this result suggests that LLMs, despite having 100 to 1000 times more parameters, are substantially more difficult to prune directly.
In this work, we address this challenge by introducing a straightforward and effective approach, termed Wanda (Pruning by Weights and activations). This technique successfully prunes LLMs to high degrees of sparsity without any need for modifying the remaining weights. This is inspired by insights from a recent study Dettmers et al. (2022) observing that a small subset of hidden state features are exceptionally large in magnitude, a property unique to LLMs. We find that augmenting the standard weight magnitude pruning metric with the input activations, is surprisingly effective as a measure for evaluating the weight importance. Specifically, we introduce a novel pruning metric, where each weight is evaluated by the product of its magnitude and the norm of the corresponding input activations, estimated using a small set of calibration data. Our method uses this metric to induce sparsity in pretrained LLMs by comparing weights locally within each output of linear layers and removing lower priority weights. Our approach is computationally efficient, able to be executed in a single forward pass, and requires minimal memory overhead.
We empirically evaluate Wanda on LLaMA (Touvron et al., 2023), one of the most effective open-sourced LLM families. Our results demonstrate Wanda can find efficient sparse networks directly from pretrained LLMs, without any retraining or weight update. Wanda outperforms magnitude pruning by a large margin and also outperforms or matches the performance of a recently proposed method for pruning LLMs (Frantar and Alistarh, 2023), while requiring a lower computational cost. We hope our work serves as a baseline for future work in this area, and encourage further exploration in understanding sparsity in LLMs.
2 Preliminaries
Magnitude Pruning (Han et al., 2015, 2016) is a standard pruning technique to induce sparsity in neural networks. Different from structured pruning approaches Xia et al. (2022b); Liu et al. (2017), magnitude pruning removes individual weights based on their magnitudes, where weights with magnitudes below a certain threshold are removed. In practice, this threshold is typically determined by comparing weights globally Liu et al. (2019) or layer-wise Zhu and Gupta (2017). Despite its simplicity, magnitude pruning has been used to find extremely sparse networks Jonathan and Michael (2019); Zhu and Gupta (2017) and now stands out as a strong baseline approach Gale et al. (2019); Blalock et al. (2020); Hoefler et al. (2021) for neural network sparsification.
Emergent Large Magnitude Features is a property specific to Transformer Vaswani et al. (2017) based large language models. Once models reach a certain scale (in practice, around 6B parameters), a small set of hidden state features emerges with significantly larger magnitudes than the remaining ones. The existence of these outlier features in large language models is demonstrated in Dettmers et al. (2022). These outlier features exhibit several intriguing characteristics. First, they have very large magnitudes, about 100 times larger than typical hidden state values (Xiao et al., 2023). Second, they are usually sparse and exist in certain feature dimensions. Finally, these outlier features are essential for the predictive capability of LLMs: zeroing out these features at inference time results in significant degradation of language modeling performance (Dettmers et al., 2022).
3 Pruning by Weights and Activations
In this section, we motivate and describe our proposed pruning approach, Wanda (Pruning by Weights and activations), which consists of two simple but essential components. First, we propose a novel pruning metric that incorporates both weights and input activations into the computation of weight importance. Second, we compare weights on a per-output basis instead of across the whole layer, a factor that we find is crucial for pruning LLMs effectively. In Figure 1, we show an overview of our pruning approach.
We first start with a motivating example. Consider a neuron with two inputs and corresponding weights: , where . Now suppose the goal is to select one weight for removal while incurring less change on the output. The standard approach of magnitude pruning would always remove weight , which may be a good strategy if input features and have similar magnitudes. However, as recently observed in LLMs Dettmers et al. (2022), the two input features can differ significantly in scale. For instance, it is possible that , and as a result, . In this case, we should remove weight instead, because this clearly exerts a smaller influence on the neuron output than removing weight .
This motivating example with the simplest linear layer hints at a major limitation of magnitude pruning: it does not take into account input activations, which could play an equally important role as weight magnitudes in determining the neuron output. For pruning LLMs, this is especially critical considering the emergent large magnitude features found within them. Thus, as the first part of our method, we propose a pruning metric designed explicitly for LLMs to handle such a limitation, while also maintaining the simplicity of magnitude pruning.
Pruning Metric. Consider a fully connected layer with weight of shape . For language models, this linear layer takes in input activations with a shape of , where and are batch and sequence dimensions respectively. For each individual weight, we propose to evaluate its importance by the product of its magnitude and the corresponding input feature norm. Specifically, the score for the current weight is defined by:
where represents the absolute value operator, evaluates the norm of th features aggregated across different tokens, and the final score is computed by the product of these two scalar values. We find that norm tends to work better than other norm functions in measuring activation magnitudes, for instance, and norm. We hypothesize this is likely because is generally a smoother metric Cortes et al. (2009); Lewkowycz and Gur-Ari (2020).
This metric is interesting in several aspects. First, when the input channel of the considered weight has large magnitude features, the weight itself tends to be assigned a larger importance score despite potentially having a low magnitude. This tackles the problem we encounter in the motivating example. The effect can be seen in Figure 1, where weights corresponding to the large magnitude feature are more likely to be preserved with our approach Wanda. Second, its computation is straightforward. Once we obtain the norm vector of input feature activations, the weight importance can be calculated using an element-wise dot product. Last, empirical evidence from our experiments suggests that this metric is robust and can be easily estimated using a modest number of calibration samples (Section 4.5), without access to the original training data.
Pruning Granularity. To motivate the second aspect of our approach, we argue that in addition to the pruning metric, the pruning granularity, i.e. the set of weights among which to compare the importance, plays a significant role in pruning LLMs. Existing pruning methods typically employ a layer-wise Zhu and Gupta (2017) or global Liu et al. (2023); Sanh et al. (2020) pruning strategy, meaning that weights are compared either within each layer or across all layers in the network. This practice has been predominant in magnitude pruning approaches, often successfully leading to highly sparse networks Jonathan and Michael (2019); Gale et al. (2019).
While the choice of pruning granularity has often been taken for granted in prior pruning approaches Jonathan and Michael (2019); Liu et al. (2023), we suggest that pruning LLMs could benefit from more local granularity levels. In our approach, we compare and remove weights on a per-output basis (per row in Figure 1), where weight importance scores are compared locally within each output neuron. Specifically, for a weight that connects input to output inside the linear layer, the group of comparison for this weight is defined as all weights connecting to output :
Under this granularity, for a pre-defined sparsity ratio , we eliminate of the weights connected to each output. This practice may seem counter-intuitive, since we are basically pruning under a stricter sparsity pattern. However, we find that it is consistently better than layer-wise pruning for LLMs. A possible reason is that removing weights at a uniform ratio per output helps prevent imbalanced pruning across different output features.
We find that for LLMs, removing weights at a per-output level not only enhances our proposed pruning metric but also the standard magnitude metric. To see if this holds true in general, we conduct additional experiments on pruning image classifiers. We observe no similar trend in image classification models, suggesting that observations regarding pruning per output on LLMs might be due to certain properties of the language modeling task. We hope this intriguing observation encourages practitioners to be more cautious in choosing the pruning granularity.
Procedure. Wanda can be implemented and integrated seamlessly within a single forward pass of the LLM model, where feature norm statistics are estimated with a set of calibration data. We provide the PyTorch code of our approach in Algorithm 1. Given a pretrained LLM, we compute our pruning metric from the initial to the final layers of the network. After pruning a preceding layer, the subsequent layer receives updated input activations, based on which its pruning metric will be computed. A recent method for pruning LLMs, SparseGPT (Frantar and Alistarh, 2023), requires iterative weight update with a sophisticated procedure in the gradual pruning process, while Wanda requires no weight update. The sparse LLM after pruning is ready to use without further training or weight adjustment.
Structured N:M Sparsity. While our method Wanda so far has been developed for unstructured sparsity, it can be easily extended to structured N:M sparsity (Zhou et al., 2021; Hubara et al., 2021; Pool and Yu, 2021). N:M sparsity requires that at most N out of every M contiguous weights to be non-zero. It can leverage NVIDIA’s sparse tensor cores (Mishra et al., 2021) to accelerate matrix multiplication in practice. Wanda can be naturally extended to structured N:M sparsity, where we only need to compare weights using the same metric among every M consecutive weights, for all weights connected to an output.
Remark. We discuss Wanda’s connection with a few existing works. SparseGPT (Frantar and Alistarh, 2023) formalizes the problem of pruning LLMs by solving a local layer-wise reconstruction problem, where their pruning metric and weight update procedure is inspired from Optimal Brain Surgeon (OBS) (Hassibi et al., 1993). The pruning metric in SparseGPT is:
Here in the denominator is the Hessian for the layer-wise reconstruction problem and is the Hessian dampening factor to avoid the collapse of inverse computation. With careful inspection, we observe that our metric in Equation 1 is similar to the above when is 0 and only the diagonal elements of the Hessian matrix are retained. This is because the diagonal of is , and the denominator in Equation 3 is now simplified to . The resulting metric is the square of our proposed metric. This simplification substantially reduces the required computation of weight importance, eliminating the need for computing any matrix inverses.
In the 1980s, LeCun et al. (LeCun et al., 1989) have set up a pioneering framework for neural network pruning named Optimal Brain Damage (OBD). It uses second-order information but ignores off-diagonal elements in Hessians for faster approximation. Later, Optimal Brain Surgeon (OBS) (Hassibi et al., 1993) develops upon OBD partly by taking into account the off-diagonal elements. Wanda can be seen as a renaissance of the pioneering work of Optimal Brain Damage (OBD) (LeCun et al., 1989) – it may be viewed as applying a process similar to OBD to each neuron, with local output reconstruction as the objective function, whereas the original OBD uses the global training objective.
| Method | Retraining | Weight Update | Calibration Data | Pruning Metric | Complexity |
| Magnitude | ✗ | ✗ | ✗ | ||
| SparseGPT | ✗ | ✓ | ✓ | ||
| Wanda (ours) | ✗ | ✗ | ✓ |
A comparison of LLM pruning methods can be found in Table 1. Here the time complexity of computing the Wanda pruning metric is reduced compared with SparseGPT because it does not involve inverse computation. Overall, our method Wanda (Pruning by Weights and activations) has several attractive properties as an approach for pruning LLMs:
1.
It maintains the simplicity of magnitude pruning in the pre-LLM era, requiring no gradient computation via back-propagation or any second-order Hessian inverses, but is also highly effective in discovering sparse networks in pretrained LLMs.
2.
Wanda can be done with a single forward pass of the LLM. At each layer, the pruned weights can be decided in one shot without an iterative procedure. In practice, Wanda can be 300 times faster in pruning LLMs compared with SparseGPT (Frantar and Alistarh, 2023).
3.
Unlike SparseGPT, our approach entails no weight updates on pruned networks, suggesting that LLMs have effective sparse sub-networks that are exact, instead of them merely existing in the neighborhood of the original weights.
4 Experiments
4.1 Setup
Models. We evaluate Wanda on the LLaMA Touvron et al. (2023) model family, a series of Transformer language models at various parameter levels, often referred to as LLaMA-7B/13B/30B/65B. We apply our pruning method to all four LLaMA models. We note that our approach is not limited to LLaMA, but is applicable to any architectures that consist of many linear layers, including other Transformer-based LLMs. We provide results for other LLM families in Appendix B.
Evaluation. We first measure the performance of pruned networks by their language modeling ability: perplexity computed on a held-out validation set. Following previous works on LLM compression Xiao et al. (2023); Frantar and Alistarh (2023); Frantar et al. (2023), we report the perplexity metric on WikiText Merity et al. (2016) validation set. While perplexity has been shown to be a stable and robust metric for language models Xia et al. (2022a), we also evaluate their zero-shot ability. We use the public evaluation benchmark EleutherAI LM Harness Gao et al. (2021). For zero-shot performance, we evaluate on seven common sense benchmarks: BoolQ Clark et al. (2019), RTE Wang et al. (2018), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2019), ARC Easy and Challenge Clark et al. (2018), and OpenbookQA Mihaylov et al. (2018). We report the accuracy on each benchmark as well as the overall mean accuracy.
Baselines. We compare our method Wanda to two prior pruning methods:
Magnitude pruning (Han et al., 2015) is a simple and strong baseline in which weights are discarded based on their magnitudes. We follow the previous practice of magnitude pruning on Transformers Gale et al. (2019); Sanh et al. (2020), where weights are compared locally among the current linear layer. Magnitude pruning has been demonstrated to outperform many other pruning methods (Gale et al., 2019; Blalock et al., 2020).
SparseGPT (Frantar and Alistarh, 2023) is a second-order pruning method based on solving a layer-wise reconstruction problem. To scale existing second-order based approaches (Frantar et al., 2022) to LLMs, an efficient weight update procedure was proposed that iterates between weight removal and weight update at each layer. To our knowledge, it is the only pruning method so far that has been demonstrated to work with LLMs.
Both Wanda and SparseGPT require calibration data to estimate certain input statistics (see Table 1). To control this variable factor, we use the exact same set of calibration data as SparseGPT, which consists of 128 sequences (2048 tokens each) sampled from the first shard of the C4 training data Raffel et al. (2020).
Sparsity. For all pruning methods, we follow the setup of SparseGPT (Frantar and Alistarh, 2023), where a uniform sparsity is imposed for all layers and there is no subsequent retraining. We skip the first embedding layer and the final classification head, as is common in pruning Transformers Sanh et al. (2020); Frantar and Alistarh (2023). Our primary approach to induce sparsity is through unstructured pruning. Considering the potential need for practical speedup, we also conduct evaluations on structured N:M sparsity Zhou et al. (2021); Hubara et al. (2021). Specifically, we provide comparisons on 4:8 and 2:4 sparsity patterns. The magnitude pruning baseline is extended to structured N:M sparsity in a similar spirit to our method which is described in the previous section.
| LLaMA | ||||||
| Method | Weight Update | Sparsity | 7B | 13B | 30B | 65B |
| Dense | - | 0 | 5.68 | 5.09 | 4.77 | 3.56 |
| Magnitude | ✗ | 50 | 17.29 | 20.21 | 7.54 | 5.90 |
| SparseGPT | ✓ | 50 | 7.22 | 6.21 | 5.31 | 4.57 |
| Wanda | ✗ | 50 | 7.26 | 6.15 | 5.24 | 4.57 |
| Magnitude | ✗ | 4:8 | 16.84 | 13.84 | 7.62 | 6.36 |
| SparseGPT | ✓ | 4:8 | 8.61 | 7.40 | 6.17 | 5.38 |
| Wanda | ✗ | 4:8 | 8.57 | 7.40 | 5.97 | 5.30 |
| Magnitude | ✗ | 2:4 | 42.13 | 18.37 | 9.10 | 7.11 |
| SparseGPT | ✓ | 2:4 | 11.00 | 9.11 | 7.16 | 6.28 |
| Wanda | ✗ | 2:4 | 11.53 | 9.58 | 6.90 | 6.25 |
4.2 Language Modeling
Unstructured Sparsity. For each of the LLaMA models, we prune it to unstructured 50% sparsity for all methods. Results are shown in Table 2. Without any weight update, Wanda outperforms the established pruning approach of magnitude pruning by a large margin. For instance, for LLaMA-7B, Wanda is able to find sparse networks with a perplexity of 7.26, significantly better than the magnitude pruning baseline 17.29. Our method also performs on par with or in most cases better than the prior reconstruction-based method SparseGPT.
Structured N:M Sparsity. We now turn our eyes to structured N:M sparsity. Note that N:M sparsity by definition is a more restrictive sparsity pattern than unstructured sparsity. Thus, it is expected that such structured sparsity patterns would lead to worse results. Results for structured 4:8 and 2:4 sparsity are shown in the lower parts of Table 2. We can see that Wanda can be easily generalized to structured N:M sparsity. Across 4:8 and 2:4 sparsity, our method consistently finds highly effective sparse sub-networks, outperforming baseline approaches in most cases, especially for larger models (e.g. 30B and 65B).
Remark. Note that in certain cases, Wanda obtains results slightly better than but somewhat close to those of SparseGPT. This may be related to the fact that our pruning metric shares an implicit connection to the OBS reconstruction error in Equation 3, as identified and explained in Section 3. While perplexity improvements in these cases over SparseGPT may not be considered substantial, Wanda is much simpler and more computationally efficient.
Varying Sparsity Levels. We conduct experiments with varying levels of sparsity for unstructured pruning, the results of which are depicted in Figure 2. It can be seen that Wanda and SparseGPT show similar trends of perplexity increase as the sparsity level gets higher. However, magnitude pruning displays a considerably more severe degradation trend.
4.3 Zero-Shot Tasks
We evaluate pruned models on downstream zero-shot tasks via prompting. Results are summarized in Table 3, where models are pruned to unstructured 50% sparsity. Averaging the accuracy over the 7 tasks under consideration, the Wanda method is able to identify effective pruned networks, showing competitiveness with baseline methods. Note that as the model gets larger in size, the accuracy drop compared to the original dense model keeps getting smaller. For task-wise performance, we observe that there are certain tasks where our approach Wanda gives consistently better results across all LLaMA models, i.e. HellaSwag, ARC-c and OpenbookQA, while among the remaining tasks, there is not a fixed superior between the two methods.
| Params | Method | BoolQ | RTE | HellaSwag | WinoGrande | ARC-e | ARC-c | OBQA | Mean |
| 7B | Dense | 71.7 | 53.4 | 58.3 | 68.0 | 67.7 | 38.6 | 28.0 | 55.1 |
| Magnitude | 53.5 | 56.3 | 44.6 | 54.5 | 53.7 | 32.3 | 21.6 | 45.2 | |
| SparseGPT | 71.5 | 56.8 | 52.7 | 65.8 | 64.3 | 36.0 | 25.0 | 53.2 | |
| Wanda | 70.3 | 53.3 | 52.9 | 64.7 | 64.1 | 37.0 | 26.4 | 52.7 | |
| 13B | Dense | 68.3 | 65.3 | 60.8 | 70.0 | 73.6 | 44.0 | 30.6 | 58.9 |
| Magnitude | 62.1 | 45.8 | 44.3 | 58.9 | 49.8 | 33.1 | 27.4 | 45.9 | |
| SparseGPT | 66.7 | 52.0 | 55.1 | 70.9 | 66.9 | 39.2 | 26.0 | 53.8 | |
| Wanda | 67.1 | 61.0 | 56.6 | 71.7 | 68.4 | 41.7 | 28.0 | 56.4 | |
| 30B | Dense | 66.9 | 61.4 | 64.8 | 72.4 | 75.3 | 46.9 | 29.4 | 59.6 |
| Magnitude | 66.6 | 53.4 | 49.7 | 63.4 | 68.8 | 40.0 | 26.2 | 52.6 | |
| SparseGPT | 71.0 | 61.4 | 61.1 | 72.0 | 73.9 | 46.0 | 31.2 | 59.5 | |
| Wanda | 70.3 | 66.8 | 62.3 | 71.1 | 74.8 | 46.5 | 32.4 | 60.6 | |
| 65B | Dense | 81.8 | 71.8 | 65.2 | 76.9 | 75.4 | 47.2 | 36.4 | 65.0 |
| Magnitude | 79.8 | 66.8 | 62.5 | 70.3 | 70.4 | 46.7 | 31.4 | 61.1 | |
| SparseGPT | 81.2 | 70.4 | 62.5 | 74.1 | 74.8 | 44.7 | 32.2 | 62.8 | |
| Wanda | 82.2 | 69.1 | 64.5 | 73.9 | 74.6 | 45.9 | 32.8 | 63.3 |
4.4 Pruning Speed
The theoretical computational complexity of Wanda is lower than SparseGPT (Table 1). Here we compare their empirical pruning speed. We measure the total pruning time only (excluding the forward pass process shared by both methods) on NVIDIA A6000 GPUs. We present the results in Table 4. Wanda incurs negligible time overhead compared with SparseGPT. The fast speed is particularly useful when pruning needs to be performed on a real-time basis, e.g., processing different samples with different sparsity levels.
| LLaMA | ||||
| Method | 7B | 13B | 30B | 65B |
| SparseGPT | 203.1 | 339.0 | 810.3 | 1353.4 |
| Wanda | 0.54 | 0.91 | 2.9 | 5.6 |
4.5 Ablation Study
We study three aspects of Wanda, aiming to answer the following questions: How does the choice of pruning configurations, i.e. pruning metric and pruning granularity, impact the performance of pruned models? How robust is our pruning method to calibration samples? Can Wanda be integrated with the OBS based weight update as used in SparseGPT? For the following ablation studies, we prune the LLaMA-7B model to unstructured 50% sparsity and report the perplexity on WikiText.
Pruning Configuration. Wanda differs from previous methods in both the pruning metric and the pruning granularity. We conduct experiments on a variety of pruning metrics and grnularities to ablate their effectiveness. The three pruning metrics can be found in Table 1. SparseGPT adopts a local granularity level inside a layer, where weights connected to 128 consecutive input channels form a group. Wanda groups weights connected with a single output channel. Therefore, we ablate two blocksize options (128 and 1) and the input/output choice. For simplicity, we use (input/output, blocksize) to denote the local pruning granularity level, e.g., (input, 1). For this experiment, we do not perform weight update for SparseGPT to focus on metric and granularity.
| Pruning Granularity | |||||
| Pruning Metric | layer | (input, 1) | (input, 128) | (output, 1) | (output, 128) |
| Magnitude: | 17.29 | 8.86 | 16.82 | 13.41 | 17.47 |
| SparseGPT: | 7.91 | 8.86 | 8.02 | 7.41 | 7.74 |
| Wanda: | 7.95 | 8.86 | 8.12 | 7.26 | 7.71 |
The results are shown in Table 5. Wanda’s default configuration delivers the best pruned model (perplexity 7.26). Interestingly, for the magnitude metric, comparing weights of the same input neuron (input, 1) yields a perplexity of 8.86, significantly better than other granularity options. Three methods also produce equivalent pruning results as under this granularity – the input is the same, thus weight ranking only depends on weight magnitude. This finding further highlights the importance of using a proper pruning granularity for pruning LLMs, even for the classical magnitude pruning.
Robustness to Calibration Samples. Our method relies on a set of calibration data to estimate input statistics of each layer. It is thus important to understand how the construction of calibration data affects the final pruned networks. We sample 5 sets of calibration data with the same size (128) from different shards of C4 training data. We find that pruned networks obtained from 5 calibration sets have a perplexity of 7.21, 7.29, 7.23, 7.25, 7.25, respectively. The average perplexity is 7.25, close to the 7.26 we reported previously. This indicates the performance does not depend too heavily on the exact selection of calibration data.
We then vary the number of calibration samples by selecting different sample sizes ranging between 1 and 256. Results are summarized in Figure 3. We see a clear difference in trend as the size of calibration data changes, where Wanda is much more robust when the calibration samples are few. Notably, even with a single sample, pruned networks obtained by our approach have a perplexity of 7.66. This may be because input norm statistics could be much easier to estimate than the full inverse hessian of the local layer-wise reconstruction problem.
Weight Update. We test if we could obtain better pruned models by incorporating the OBS weight update process proposed in SparseGPT Frantar and Alistarh (2023). We apply the OBS weight update procedure used in SparseGPT on the kept weights produced by both magnitude pruning and our approach Wanda. We evaluate the perplexity on both the calibration data and the validation data from WikiText.
Results are summarized in Table 6. The result on magnitude pruning does validate the effectiveness of the weight update procedure, where it improves perplexity significantly on both data splits. However, when the pruned model is obtained by our method Wanda, it no longer brings improvement but leads to an increase of perplexity (from 7.26 to 7.32) on validation data, suggesting that in this case the updated model may overfit to the calibration samples.
5 Analysis
In this section, we first study whether Wanda can be a general pruning approach beyond LLMs. Next, we show exploratory results on using parameter efficient fine-tuning techniques to recover performance drop during the pruning procedure.
5.1 Generality beyond Pruning LLMs
While the main focus of this work is on pruning LLMs, the surprising effectiveness of Wanda leads to a clear question: how would Wanda perform against magnitude pruning on tasks where the latter has been widely used? We thus conduct a study on ImageNet-1K (Deng et al., 2009), a standard image classification task where magnitude pruning has been extensively studied and accepted as a strong baseline Gale et al. (2019); Blalock et al. (2020).
We consider two modern vision architectures: ConvNeXt Liu et al. (2022) and Vision Transformer (ViT) Dosovitskiy et al. (2021). We choose these two architectures mainly for two reasons: first, as LLMs are based on Transformers Vaswani et al. (2017), we would like to test if our observations on LLMs so far still hold on Transformers for other tasks; second, as we are evaluating on image classification, we are interested in examining how they work with ConvNet models, with ConvNeXt being a representative architecture.
We use two ImageNet-1K pretrained models: ConvNeXt-B and ViT-B, which have a top-1 accuracy of 83.8% and 81.8% Touvron et al. (2021) respectively. We prune the linear layers only (for ConvNeXt, this includes equivalent 11 convolution layers). We find that for both models, pruning uniformly on a per layer basis is consistently better than a per output basis (detailed results are shown in Appendix B). We thus compare our proposed metric and the standard magnitude metric by pruning on a per layer basis. Results are shown in Figure 4. Our novel pruning metric leads to better results than magnitude pruning, especially at high sparsities (e.g. 70%, 80%). However, we find differences in the accuracy of pruned networks can easily be mitigated via retraining for a few epochs, indicating that Wanda might be more suitable for cases where retraining is not computationally feasible, such as in LLMs.
5.2 LoRA Fine-tuning
We explore using parameter efficient fine-tuning (PEFT) techniques to recover performance of pruned LLM models. We use a popular PEFT method LoRA Hu et al. (2021), which has been widely adopted for task specific fine-tuning of LLMs. However, here we are interested in recovering the performance loss of LLMs during pruning, thus we perform a more general “fine-tuning” where the pruned networks are trained with an autoregressive objective on C4 dataset. We enforce a limited computational budget (1 GPU and 5 hours). We find that we are able to restore performance of pruned LLaMA-7B (unstructured 50% sparsity) with a non-trivial amount, reducing zero-shot WikiText perplexity from 7.26 to 6.87. The additional parameters introduced by LoRA is only 0.06%, leaving the total sparsity level still at around 50% level.
6 Related Work
Network Pruning and Sparsity. Pruning is a popular technique for compressing neural networks through the elimination of weights, yielding sparse networks (LeCun et al., 1989; Hassibi et al., 1993; Han et al., 2015; Liu et al., 2019; Zhu et al., 2018). Pruning methods can be broadly categorized into structured and unstructured approaches. Structured pruning methods Xia et al. (2022b); Liu et al. (2017); Molchanov et al. (2017); Fan et al. (2020) remove entire structured components of a network, facilitating efficient GPU speedups, while unstructured methods like magnitude pruning operate at the individual weight level, maintaining performance even at higher sparsity levels. Existing pruning methods usually require either modifications to the training procedure Louizos et al. (2018); Sanh et al. (2020); Gale et al. (2019); Kusupati et al. (2020), retraining the pruned networks to regain accuracy Han et al. (2015, 2016); Liu et al. (2019), or an even more computationally intensive iterative retraining process Jonathan and Michael (2019); Zhu and Gupta (2017); Renda et al. (2020); Frankle et al. (2020). However, scaling these methods to LLMs with billions of parameters presents a challenge, as the required training process demands substantial computational resources Touvron et al. (2023); Zhang et al. (2022).
Pruning with Limited Data. Most related to our approach is a recent line of work on pruning with limited data Hubara et al. (2021); Frantar et al. (2022); Frantar and Alistarh (2022); Kwon et al. (2022). Such methods require no modification to the original training procedure and also no retraining of the pruned networks on the full training dataset. The primary aim of these methods is to preserve performance during the pruning procedure, assuming access to a limited and small amount of data, also referred to as the calibration data. In order to mitigate the accuracy drop, a layer-wise reconstruction problem (Hubara et al., 2021) is solved to minimize the change of output evaluated on the calibration data. Existing popular solvers Frantar et al. (2022); Singh and Alistarh (2020) for the layer-wise reconstruction problem rely on heavy computation of second-order Hessian inverses, which do not scale to the large hidden state size of LLMs. SparseGPT (Frantar and Alistarh, 2023) develops an efficient weight update procedure for LLMs via synchronized second-order Hessian updates.
Emergent Properties of LLMs. Our work is also related to recent studies on the existence of large magnitude outlier features in Transformer-based language models (Timkey and Schijndel, 2021; Bondarenko et al., 2021; Xiuying et al., 2022; Luo et al., 2021; Puccetti et al., 2022). Dettmers et al. (2022) demonstrates that when LLMs exceed a certain parameter scale (e.g. 6B), large magnitude features start to emerge and strongly affect all layers, which is considered as one of the emergent properties of LLMs (Dettmers et al., 2022; Wei et al., 2022a; Schaeffer et al., 2023). They further identify these emergent features as the cause of existing quantization methods’ failures. This observation has spurred the development of various quantization schemes Dettmers et al. (2022); Xiao et al. (2023); Lin et al. (2023); Dettmers et al. (2023) tailored specifically for LLMs to manage these outlier features. Our work extends this understanding, demonstrating that these emergent large magnitude features should also serve as pivotal indicators for determining which weights to prune in LLMs.
7 Conclusion
In this work, we propose a simple and effective method for pruning Large Language Models (LLMs). Inspired by the recent discovery of emergent large features in LLMs, our approach, termed Wanda (Pruning by Weights and activations), removes weights with the smallest magnitudes multiplied by the corresponding input activation norms, on a per-output basis. Without the need for any retraining or weight update procedures, Wanda is able to identify effective sparse networks within pretrained LLMs. Wanda could be used as a general pruning approach beyond LLMs. We hope our work contributes to a better understanding of sparsity in LLMs.
Acknowledgments. We thank Yonghao Zhuang for valuable discussions. Mingjie Sun and Anna Bair were supported by funding from the Bosch Center for Artificial Intelligence.
Appendices
Appendix A Experimental Details
Our experiments are conducted on NVIDIA RTX A6000 GPUs with 50 GB memory. For pruning LLaMA models, we first load these models onto GPUs in 16-bit floating-point format. All subsequent procedures (e.g. pruning) are conducted directly on GPUs. For zero-shot tasks, we use the evaluation framework from https://github.com/EleutherAI/lm-evaluation-harness/.
Pruning Image Classifiers. We use the pretrained ConvNeXt-B model available at https://github.com/facebookresearch/ConvNeXt. For ViT, we use the DeiT-B model from https://github.com/facebookresearch/deit. The ConvNeXt-B and DeiT-B classifiers have 89M and 86M parameters respectively. As described in Section 5, we focus on pruning linear layers, which comprise the majority of the parameters in these two architectures. For ConvNeXt, the 11 convolutions in the its blocks can be viewed as linear layers, which account for approximately 95% of the total model parameters. For calibration data on ImageNet, we sample 4096 images from the training set, using them as the calibration set for our entire analysis. We find 4096 samples leads to a stable result for our pruning metric, beyond which we notice only marginal effect.
LoRA Fine-tuning. We adopt the LoRA fine-tuning code on LLaMA available at https://github.com/tloen/alpaca-lora. Our fine-tuning objective is the pretraining autoregressive language modeling objective. For the LoRA adapter, we update two weight matrices, and , in the self-attention module with a rank of 8. The fine-tuning process is conducted on the C4 training set, operating within a restricted computational budget that involves a single A6000 GPU and a time constraint of 5 hours.
Appendix B Additional Results
B.1 Pythia and OPT
While prior public LLM models may not be as powerful as LLaMA, we test the generality of our approach by evaluating it on two representative LLM families before the release of LLaMA, namely Pythia [Biderman et al., 2023] and OPT Zhang et al. [2022]. We select two models with approximately the same level of total parameters: Pythia-12b 111https://huggingface.co/EleutherAI/pythia-12b and OPT-13b 222https://huggingface.co/facebook/opt-13b. The results (unstructured sparsity) are shown in Table 7. We can see that magnitude pruning deteriorates significantly faster on Pythia-12b and OPT-13b in comparison to LLaMA models. For instance, even a sparsity ratio of 20% pushes the perplexity of pruned models to a meaningless level (greater than 1000). In contrast, our method Wanda consistently finds exact and effective sparse networks within both pretrained LLMs.
| Sparsity | ||||||||
| Model | Dense | Pruning Method | Weight Update | 10% | 20% | 30% | 40% | 50% |
| Pythia-12b | 8.59 | Magnitude | ✗ | 127.76 | 2e5 | 7e5 | 2e5 | 3e5 |
| SparseGPT | ✓ | 8.59 | 8.65 | 8.86 | 9.39 | 11.02 | ||
| Wanda | ✗ | 8.59 | 8.60 | 8.85 | 9.31 | 11.27 | ||
| OPT-13b | 10.13 | Magnitude | ✗ | 14.45 | 9e3 | 1e4 | 1e4 | 1e4 |
| SparseGPT | ✓ | 10.11 | 10.10 | 10.12 | 10.35 | 11.19 | ||
| Wanda | ✗ | 10.09 | 10.07 | 10.09 | 10.63 | 11.42 | ||
B.2 Image Classifiers
In Figure 5, we present detailed results on pruning image classifiers. We report the accuracy of pruned models without any subsequent retraining or fine-tuning. We can see that for both ConvNeXt and DeiT, layer-wise pruning is slightly better than pruning per output.
In addition, we extend our analysis to a ViT model from the original paper Dosovitskiy et al. [2021]: ViT-L trained on ImageNet-1K with a top-1 accuracy 78.9%. To differentiate from DeiT, we refer to this pretrained model as original ViT. Results are shown in Figure 6. Interestingly, in this case, we observe that pruning on a per output basis is superior to layerwise pruning, implying that the choice of pruning granularity may be different for different models.
Notably, we also find that the accuracy differences between Wanda and magnitude pruning in pruning image classifiers can be effectively mitigated by fine-tuning for a few epochs.